# Optimized mechanical quadrature squeezing beyond the 3 dB limit via gradient-descent algorithm

### v2

@@ III GENERATION OF MECHANICAL QUADRATURE SQUEEZING
Figure 5: (a-c) The squeezing degree $\mathcal{S}_{b}$ versus the scaled evolution time $\omega_{\mathrm{m}}t$ under different sideband parameters $\kappa/\omega_{\mathrm{m}}=0.5$, 1.0, and 1.5. The inset in panel (b) exhibits the evolution of $\mathcal{S}_{b}$ for the time durations from $115.6$ to $120$. (d-f) Evolution of driving amplitude and (g-i) phase corresponding to panels (a-c). (j) The dynamic evolution of the mean phonon number under different sideband-resolution parameters in panels (a-c). Other parameters are consistent with those given in Fig. 1.
IV DISCUSSIONS AND CONCLUSION
@@
In this section, we present some discussions on the physical mechanism and experimental implementation of this scheme and conclude this work. The mechanical squeezing generation can be understood by analyzing the variance of the mechanical rotating-quadrature operator $\Delta X_{b}^{2}(\theta,t)=\mathbf{V}_{33}(t)\cos^{2}\theta+\mathbf{V}_{44}(t)\sin^{2}\theta+\frac{1}{2}[\mathbf{V}_{34}(t)+\mathbf{V}_{43}(t)]\sin(2\theta)$. To be more clear, we re-express the variance as
@@
Since the mechanical squeezing occurs when $\Delta X_{b}^{2}(\theta,t)<1/2$, our goal is to realize $\langle b^{\dagger}b\rangle+\mathrm{Re}[\langle b^{\dagger}b^{\dagger}\rangle]\cos(2\theta)-2\mathrm{Im}[\langle b^{\dagger}b^{\dagger}\rangle]\sin(2\theta)<0$, which is feasible because the second-order moment $\langle b^{\dagger}b^{\dagger}\rangle$ could be a complex number. Meanwhile, the value of $\langle b^{\dagger}b\rangle$ must be a positive number, and then the reduction of $\langle b^{\dagger}b\rangle$ is beneficial to the squeezing generation, which explains why the mechanical resonator is cooled [Figs. 2(j), 2(k), 4(e), and 5(j)]. In addition, under the rotating-wave approximation, we have two sets of equations governing all these second-order moments. The first set includes $\langle a^{\dagger}a\rangle$, $\langle b^{\dagger}b\rangle$, $\langle a^{\dagger}b\rangle$, and $\langle ab^{\dagger}\rangle$, while the second set includes $\langle a^{\dagger}a^{\dagger}\rangle$, $\langle aa\rangle$, $\langle b^{\dagger}b^{\dagger}\rangle$, $\langle bb\rangle$, $\langle a^{\dagger}b^{\dagger}\rangle$, and $\langle ab\rangle$. Initially, with the condition $\mathbf{X}(0)=[0,\bar{n}_{\mathrm{m}},0,0,0,0,0,0,0,0]^{\mathrm{T}}$, the value of $\langle b^{\dagger}b^{\dagger}\rangle$ remains constantly zero, making mechanical squeezing unattainable regardless of how we optimize the driving field. In the presence of the counter-rotating terms, the equations of motion for these ten second-order moments are coupled to each other. Consequently, $\langle b^{\dagger}b^{\dagger}\rangle$ is no longer zero. In this case, we can obtain $\langle b^{\dagger}b\rangle+\mathrm{Re}[\langle b^{\dagger}b^{\dagger}\rangle]\cos(2\theta)-2\mathrm{Im}[\langle b^{\dagger}b^{\dagger}\rangle]\sin(2\theta)<0$ via optimizing the pulsed driving. These analyses indicate that the counter-rotating terms play a crucial role in the generation of mechanical quadrature squeezing.
@@
We now present some discussions on the experimental feasibility of this scheme. The system under considerations is a general optomechanical system, which has been implemented in various optomechanical platforms, such as optical microresonators ParkNP2009 ; SchliesserNP2009 ; VerhagenNature2012 , electromechanical systems HertzbergNP2010 ; MasselN2011 ; MasselNC2012 , photonic crystal nanobeams TeufelNature2011 ; Eichenfield2009 ; ChanNT2009 , and Fabry-Pérot cavities HammererNature2009 , in which the cavities can be driven by tunable pulsed fields. We considered the linearized optomechanical interaction, which has been widely demonstrated in cavity optomechanical systems. Moreover, we use the experimentally accessible parameters in our numerical simulations. Concretely, the used parameters are $g_{0}/\omega_{\mathrm{m}}=4\times 10^{-5}$, $\kappa/\omega_{\mathrm{m}}=0.2$, and $\gamma/\omega_{\mathrm{m}}=2\times 10^{-6}$, which have been reported in experiments VerhagenNature2012 . In addition, the pulsed driving amplitude and phase used to generate squeezing are moderate in size and continuous and smooth in shape, which confirm the experimental realization of the pulsed driving.
@@
In conclusion, we have presented a scheme for generating mechanical quadrature squeezing in a typical optomechanical system via gradient-descent algorithm. The generated mechanical squeezing can exceed the 3 dB steady-state limit and ultrafast squeezing preparation within one mechanical oscillation period can be realized. The optimal driving amplitude and phase corresponding to these generated squeezings have been presented. Our scheme will pave the way for exploiting optimal quantum control in quantum optics and quantum information science.
@@ Acknowledgements.
The authors thank Dr. Ye-Xiong Zeng and Dr. Yue-Hui Zhou for helpful discussions. J.-Q.L. was supported in part by National Natural Science Foundation of China (Grants No. 12175061, No. 12247105, and No. 11935006), the Science and Technology Innovation Program of Hunan Province (Grant No. 2021RC4029), and Hunan Provincial Major Science and Technology Program (Grant No. 2023ZJ1010).
@@ Appendix A Derivation of the variation $\delta[\Delta X_{b}^{2}(\theta,T)]/\delta\mathcal{Q}_{m}$
where $\mathcal{Q}$ could be eithor $\Omega$ or $\phi$, $m$ is the iteration number, and $\chi_{\mathcal{Q}}$ is the learning rate. To calculate Eq. (16), we need to calculate the variation of $\Delta X_{b}^{2}(\theta,T)$ with respect to the pulse amplitude $\Omega(s)$ and phase $\phi(s)$. Based on Eq. (12), we can obtain the result
$\!\frac{\delta\Delta X_{b}^{2}(\theta,T)}{\delta\mathcal{Q}(s)}=\frac{\delta\mathrm{V}_{33}(T)}{\delta\mathcal{Q}(s)}\cos^{2}\theta+\frac{\delta\mathrm{V}_{44}(T)}{\delta\mathcal{Q}(s)}\sin^{2}\theta+\left[\frac{\delta\mathrm{V}_{34}(T)}{\delta\mathcal{Q}(s)}+\frac{\delta\mathrm{V}_{43}(T)}{\delta\mathcal{Q}(s)}\right]\sin(2\theta).$ (17)
According to Eq. (17), we can further calculate the results of $\delta\mathrm{V}_{33}(T)/\delta\mathcal{Q}(s)$, $\delta\mathrm{V}_{44}(T)/\delta\mathcal{Q}(s)$, $\delta\mathrm{V}_{34}(T)/\delta\mathcal{Q}(s)$, and $\delta\mathrm{V}_{43}(T)/\delta\mathcal{Q}(s)$. The variation of the covariance-matrix elements with respect to the pulse amplitude $\Omega(t)$ and phase $\phi(t)$ can be expressed as a linear combination of the variation of these second-order moments with respect to the driving amplitude $\Omega(t)$ and phase $\phi(t)$,
@@
It can be seen from Eq. (18) that, to obtain the result in Eq. (17), we need to further calculate the variations of these second-order moments with respect to $\Omega(t)$ and $\phi(t)$. This can be achieved by taking the variation with respect to $\Omega(s)$ and $\phi(s)$ on both sides of Eq. (8), namely,
@@
Since $\mathcal{Q}(s)$ is a function of $s$, $\delta\dot{\mathbf{X}}(t)/\delta\mathcal{Q}(s)$ in Eq. (19) can be expressed as $\frac{d}{dt}\frac{\delta{\mathbf{X}}(t)}{\delta\mathcal{Q}(s)}$ Greiner1996FQ . Under the initial condition $\delta\mathbf{X}(0)/\delta\mathcal{Q}(s)=[0,...,0]_{10\times 1}^{\text{T}}$, the solution of Eq. (19) at the target time $T$ can be obtained as

### v3

@@ III GENERATION OF MECHANICAL QUADRATURE SQUEEZING
Figure 5: (a)-(c) The squeezing degree $\mathcal{S}_{b}$ vs the scaled evolution time $\omega_{\mathrm{m}}t$ under different sideband parameters $\kappa/\omega_{\mathrm{m}}=0.5$, 1.0, and 1.5. The inset in panel (b) exhibits the evolution of $\mathcal{S}_{b}$ for the time durations from $115.6$ to $120$. (d)-(f) Evolution of driving amplitude and (g)-(i) phase corresponding to panels (a)-(c). (j) The dynamic evolution of the mean phonon number under different sideband-resolution parameters in panels (a)-(c). The two insets show the enlarged view at the peak of 3.2 dB. Other parameters are consistent with those given in Fig. 1.
The ultrafast generation of mechanical quadrature squeezing within one mechanical oscillation period is a desired task from the viewpoint of transient evolution [65, 66]. Below, we investigate the generation of mechanical squeezing within this short timescale. Figure 4(a) shows the loss function $\Delta X_{b}^{2}(\pi/2,T=6\omega_{\mathrm{m}}^{-1})$ versus the iteration number related to $\mathcal{S}_{b}=1$. Here we can see that the $\Delta X_{b}^{2}(\pi/2,T=6\omega_{\mathrm{m}}^{-1})$ gradually decreases from the initial value 319 to 0.397 as the iteration number increases, verifying the validity of the gradient-descent algorithm. The dynamic evolution of the squeezing degree $\mathcal{S}_{b}$ after the last iteration ($m=6180$) is shown in Fig. 4(b). We see that the squeezing occurs in the last extremely short duration due to the harsh time condition, but it is still continuous, as shown in the inset of Fig. 4(b). The results in Figs. 4(c) and 4(d) indicate that a large driving amplitude is required to achieve $\mathcal{S}_{b}=1$ within one mechanical oscillation period and the phase has no distinct signature. We also exhibit in Fig. 4(e) the mean phonon number $\langle b^{\dagger}b\rangle$ versus the scaled evolution time at this time. Here we see that $\langle b^{\dagger}b\rangle$ will be larger than the initial value due to the larger driving amplitude, and then goes to a cooled state with dozens of mean phonons. This indicates that the creation of mechanical squeezing does not necessarily require the ground-state cooling of the mechanical resonator.
We also investigate the influence of the sideband-resolution parameter on the squeezing generation. In Figs. 5(a) to 5(c), we plot the squeezing degree $\mathcal{S}_{b}$ versus the scaled evolution time $\omega_{\mathrm{m}}t$ corresponding to different sideband-resolution parameters $\kappa/\omega_{\mathrm{m}}=0.5$, 1.0, and 1.5. We show the driving amplitude and phase as functions of $\omega_{\mathrm{m}}t$ for different sideband-resolution parameters in Figs. 5(d)-5(f) and Figs. 5(g)-5(i). To maintain $\mathcal{S}_{b}=1$, a larger driving amplitude is needed for a larger value of $\kappa/\omega_{\mathrm{m}}$, and the phase oscillation becomes more severe for a larger $\kappa/\omega_{\mathrm{m}}$. The corresponding dynamic evolution of the mean phonon number $\langle b^{\dagger}b\rangle$ is shown in Fig. 5(j). Here we can see that the mechanical resonator is cooled from the initial 100 phonons to a few phonons. The better the sideband-resolution parameter that is selected, the deeper the cooling of the mechanical resonator. Note that a large $\kappa/\omega_{\mathrm{m}}$ will result in a fast dissipation of the cavity photons, thereby affecting the optomechanical interaction and leading to the failure of squeezing generation. We mention that the thermal excitations are usually harmful to the generation and maintenance of quantum signatures. For the generation of strong quadrature squeezing, optimal parameters should be chosen such that the thermal excitations can be decreased to a near-ground-state cooling level.
Figure 6: The squeezing degree $\mathcal{S}_{b}$ versus the scaled evolution time $\omega_{\mathrm{m}}t$ in the presence of (a) driving amplitude and (b) phase noises. The driving amplitude is modified by adding a scaled random amplitude from a normal distribution with a mean of 0 and a standard deviation of 200. Similarly, the driving phase is modified by adding a scaled random phase from a normal distribution with a mean of 0 and a standard deviation of 0.2. Other parameters are the same as those in Fig. 2(c).
IV DISCUSSIONS
In this section, we present some discussions concerning the influences of the driving amplitude and phase noises on the squeezing generation, the physical mechanism of the present squeezing-generation method, the experimental implementation of this scheme, and the applications of the optimization methods in cavity optomechanics.
IV.1 Influences of the driving amplitude and phase noises on squeezing generation
In realistic situations, the driving laser inevitably possesses amplitude and phase noises, which can be understood as random small deviations on the amplitude and phase. To evaluate the influences of laser amplitude and phase noises on the squeezing-generation performance, we add Gaussian random noises to the optimized pulse driving amplitude and phase at each step. Figures 6(a) and 6(b) display the squeezing degree $\mathcal{S}_{b}$ as a function of the scaled evolution time $\omega_{\mathrm{m}}t$ after introducing the Gaussian random noises to the driving amplitude and phase, respectively. The ten solid curves represent the evolution of the squeezing degree $\mathcal{S}_{b}$ under ten distinct random conditions, while the purple dotted curve represents the average of these ten evolutions. We can see that the 11 curves almost overlap, which indicates that the optimized driving field is stable against both the amplitude and phase noises.
IV.2 Physical mechanism of the present squeezing-generation method
@@
In our scheme, the generation of mechanical squeezing can be understood by analyzing the variance of the mechanical rotating-quadrature operator $\Delta X_{b}^{2}(\theta,t)=\mathbf{V}_{33}(t)\cos^{2}\theta+\mathbf{V}_{44}(t)\sin^{2}\theta+\frac{1}{2}[\mathbf{V}_{34}(t)+\mathbf{V}_{43}(t)]\sin(2\theta)$. To clarify this point, we re-express the rotating-quadrature operator as
@@
Mechanical squeezing occurs when $\Delta X_{b}^{2}(\theta,t)<1/2$, which leads to the condition
$\langle b^{\dagger}b\rangle+\mathrm{Re}[\langle b^{\dagger}b^{\dagger}\rangle]\cos(2\theta)-2\mathrm{Im}[\langle b^{\dagger}b^{\dagger}\rangle]\sin(2\theta)<0.$ (16)
This inequality can be satisfied because the second-order moment $\langle b^{\dagger}b^{\dagger}\rangle$ could be a complex number, while $\langle b^{\dagger}b\rangle$ must be positive. Our numerical simulations indicate the mechanical cooling is accompanied by the physics process, and the mechanical resonator is cooled to a relatively small mean phonon number [Figs. 2(j), 2(k), 4(e), and 5(j)]. Therefore, the generation of the mechanical squeezing is primarily determined by the values of $\mathrm{Re}[\langle b^{\dagger}b^{\dagger}\rangle]$ and $\mathrm{Im}[\langle b^{\dagger}b^{\dagger}\rangle]$.
@@
Under the rotating-wave approximation, the second-order moments are governed by two separate sets of equations. The first set includes $\langle a^{\dagger}a\rangle$, $\langle b^{\dagger}b\rangle$, $\langle a^{\dagger}b\rangle$, and $\langle ab^{\dagger}\rangle$, while the second set includes $\langle a^{\dagger}a^{\dagger}\rangle$, $\langle aa\rangle$, $\langle b^{\dagger}b^{\dagger}\rangle$, $\langle bb\rangle$, $\langle a^{\dagger}b^{\dagger}\rangle$, and $\langle ab\rangle$. With the initial condition $\mathbf{X}(0)=[0,\bar{n}_{\mathrm{m}},0,0,0,0,0,0,0,0]^{\mathrm{T}}$, the value of $\langle b^{\dagger}b^{\dagger}\rangle$ remains constantly zero, leading to a negligible mechanical squeezing regardless of how we optimize the driving fields. The counter-rotating terms in the coupling will mix the equations of motion for these ten second-order moments. In this case, we can obtain $\langle b^{\dagger}b\rangle+\mathrm{Re}[\langle b^{\dagger}b^{\dagger}\rangle]\cos(2\theta)-2\mathrm{Im}[\langle b^{\dagger}b^{\dagger}\rangle]\sin(2\theta)<0$ via optimizing the pulsed driving, indicating the generation of squeezing. These analyses indicate that the critical role of the counter-rotating terms is the generation of mechanical quadrature squeezing.
Meanwhile, we examined the single-mode squeezing properties of the two-mode squeezed vacuum state and the state obtained by applying the two-mode squeezing operator to the direct product of vacuum and thermal states. We find that there is no single-mode quadrature squeezing in these two cases, indicating that the pure counter-rotating-wave terms are insufficient for generating single-mode squeezing. Based on these analyses, we deduce that the generation of single-mode quadrature squeezing arises from the combined effects of both rotating-wave and counter-rotating-wave terms. Since both the linearized optomechanical-coupling strength $G(t)$ and the normalized driving detuning $\Delta(t)$ depend on the evolution time, these provide the possibility to induce the physical process for effective two-phonon interactions, which is the physical origin for generating single-mode squeezing.
IV.3 Experimental implementation of the scheme
@@
We now present some discussions on the experimental feasibility of this scheme. The system under considerations is a general optomechanical system, which has been implemented in various optomechanical platforms, such as optical microresonators [67, 68, 69], electromechanical systems [70, 71, 72], photonic crystal nanobeams [5, 73, 74], and Fabry-Pérot cavities [75], in which the cavities can be driven by tunable pulsed fields. We considered the linearized optomechanical interaction, which has been widely demonstrated in cavity optomechanical systems. Moreover, we use the experimentally accessible parameters in our numerical simulations. Concretely, the used parameters are $g_{0}/\omega_{\mathrm{m}}=4\times 10^{-5}$, $\kappa/\omega_{\mathrm{m}}=0.2$, and $\gamma/\omega_{\mathrm{m}}=2\times 10^{-6}$, which have been reported in experiments [69]. In addition, the pulsed driving amplitude and phase used to generate squeezing are moderate in size and continuous and smooth in shape, which confirm the experimental realization of the pulsed driving.
IV.4 Applications of the optimization methods in
cavity optomechanics
To broaden the use of our optimization method in the field of cavity optomechanics, we investigate its applicability across two primary research regimes: the strong driving linearization regime and the single-photon strong-coupling regime.
In the strong driving linearization regime, the physical properties of the optomechanical system are governed by the covariance matrix. For this linearized optomechanical system, we can control the dynamics of the system by optimizing the driving field, which is an adjustable control parameter. For example, we can optimize the physical topics such as optomechanical cooling, entanglement, and squeezing. Moreover, this optimization problem can be extended and applied to multi-mode optomechanical systems.
In the single-photon strong-coupling regime, the optomechanical system can be described as a multilevel system in the eigen representation of the photon-number-dependent displaced oscillator system. The driving field will induce the transitions between the states associated with neighboring photon numbers. Here, the transition magnitudes depend on the external driving field, which provide the means for controlling the systems. Typically, the optimization allows for several potential applications, including the enhancement of the photon-blockade effect and the preparation of non-classical states such as mechanical number states. By implementing these optimizations, we can deepen both the understanding and the practical applications of the cavity optomechanical systems.
V CONCLUSION
@@
In conclusion, we have presented a scheme for generating mechanical quadrature squeezing in a typical optomechanical system via gradient-descent algorithm. The generated mechanical squeezing can exceed the 3-dB steady-state limit and ultrafast squeezing preparation within one mechanical oscillation period can be realized. The optimal driving amplitude and phase corresponding to these generated squeezings have been presented. Our scheme will pave the way for exploiting optimal quantum control in quantum optics and quantum information science.
@@ Acknowledgements.
The authors thank Dr. Ye-Xiong Zeng and Dr. Yue-Hui Zhou for helpful discussions. J.-Q.L. was supported in part by National Natural Science Foundation of China (Grants No. 12175061, No. 12247105, and No. 11935006), National Key Research and Development Program of China (Grant No. 2024YFE0102400), and Hunan Provincial Major Sci-Tech Program (Grant No. 2023ZJ1010). Y.-H.L. was supported in part by Hunan Provincial Postgraduate Research and Innovation project (Grant No. CX20240530).
@@ Appendix A Derivation of the variation $\delta[\Delta X_{b}^{2}(\theta,T)]/\delta\mathcal{Q}_{m}$
where $\mathcal{Q}$ could be either $\Omega$ or $\phi$, $m$ is the iteration number, and $\chi_{\mathcal{Q}}$ is the learning rate. To calculate Eq. (17), we need to calculate the variation of $\Delta X_{b}^{2}(\theta,T)$ with respect to the pulse amplitude $\Omega(s)$ and phase $\phi(s)$. Based on Eq. (12), we can obtain the result
$\small\!\frac{\delta\Delta X_{b}^{2}(\theta,T)}{\delta\mathcal{Q}(s)}\!\!$ $=$ $\!\!\frac{\delta\mathrm{V}_{33}(T)}{\delta\mathcal{Q}(s)}\cos^{2}\theta+\frac{\delta\mathrm{V}_{44}(T)}{\delta\mathcal{Q}(s)}\sin^{2}\theta$ (18)
$+\left[\frac{\delta\mathrm{V}_{34}(T)}{\delta\mathcal{Q}(s)}+\frac{\delta\mathrm{V}_{43}(T)}{\delta\mathcal{Q}(s)}\right]\sin(2\theta).$
According to Eq. (18), we can further calculate the results of $\delta\mathrm{V}_{33}(T)/\delta\mathcal{Q}(s)$, $\delta\mathrm{V}_{44}(T)/\delta\mathcal{Q}(s)$, $\delta\mathrm{V}_{34}(T)/\delta\mathcal{Q}(s)$, and $\delta\mathrm{V}_{43}(T)/\delta\mathcal{Q}(s)$. The variation of the covariance-matrix elements with respect to the pulse amplitude $\Omega(s)$ and phase $\phi(s)$ can be expressed as a linear combination of the variation of these second-order moments with respect to the driving amplitude $\Omega(s)$ and phase $\phi(s)$:
@@
It can be seen from Eq. (19) that, to obtain the result in Eq. (18), we need to further calculate the variations of these second-order moments with respect to $\Omega(s)$ and $\phi(s)$. This can be achieved by taking the variation with respect to $\Omega(s)$ and $\phi(s)$ on both sides of Eq. (8), namely,
@@
Since $\mathcal{Q}(s)$ is a function of $s$, $\delta\dot{\mathbf{X}}(t)/\delta\mathcal{Q}(s)$ in Eq. (20) can be expressed as $\frac{d}{dt}\frac{\delta{\mathbf{X}}(t)}{\delta\mathcal{Q}(s)}$ [76]. Under the initial condition $\delta\mathbf{X}(0)/\delta\mathcal{Q}(s)=[0,...,0]_{10\times 1}^{\text{T}}$, the solution of Eq. (20) at the target time $T$ can be obtained as